OpenStudy (anonymous):

What is a rule for the function identified by this set of ordered pairs?(1,1), (2,4), (3,9), (4,16), (5,25) A. double the number B. square the number C. triple the number D. multiply the number by 4 What is a rule for the function identified by this set of ordered pairs?(1,1), (2,4), (3,9), (4,16), (5,25) A. double the number B. square the number C. triple the number D. multiply the number by 4 @Mathematics

OpenStudy (anonymous):

b. square the number f(x)=x^2

OpenStudy (anonymous):

It is B. 1^2 = 1: (1,1) 2^2 = 4: (2,4) 3^2 = 9: (3,9) 4^2 = 16: (4,16) etc.